OpenStudy (anonymous):

SOMEONE PLEASE HELP Find the first six terms of the sequence a1=-8, an=5*an-1

OpenStudy (anonymous):

first term is \(-8\) second term is \(5\times (-8)\) third term is \(5\times 5\times (-8)=5^2\times (-8)\) fourth term is \(5^3\times (-8)\) two more to go

OpenStudy (anonymous):

Ohh okay thank you so much

OpenStudy (anonymous):

yw